OpenStudy (anonymous):

A segment with endpoints F (4, 9) and G (7, 2) is divided by a point H such that FH and GH form a 1:3 ratio. Find H. (5 points) Select one: a. (5.3, 6) b. (4.75, 7.25) c. (4.1, 7.4) d. (5.25, 5.75)

OpenStudy (anonymous):

@whpalmer4 @ganeshie8

OpenStudy (anonymous):

coordinate of the point which divides the join of the points (x1,y1) and (x2,y2) in the ratio of m:n is given by (mx2+nx1)/(m+n) , (my2+ny1)/(m+n)

OpenStudy (anonymous):

so what would it be

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

(3*4 + 1*7)(1+3) ,(3*9+1*2)/(1+3) =19/4,29/4 =(4.75, 7.25)

OpenStudy (anonymous):

thnx can you help me on some more please